OpenStudy (anonymous):

solve for y y=-x+6 y=5x-6

OpenStudy (anonymous):

-x+6=5x-6 6=6x-6 12=6x x=2

OpenStudy (maheshmeghwal9):

hence,-x+6=5x-6 according to question and u

OpenStudy (maheshmeghwal9):

take variable one side and constant one side.i.e.;

OpenStudy (maheshmeghwal9):

6+6=5x+x=>12=6x

OpenStudy (maheshmeghwal9):

Now solve for x

OpenStudy (maheshmeghwal9):

x=12/6 =>x=2. ans.......

OpenStudy (maheshmeghwal9):

do u understand it now?

OpenStudy (anonymous):

yes i need y

OpenStudy (anonymous):

y=6-x y=6-2 y=4

OpenStudy (maheshmeghwal9):

So, y=-x+6=-2+4=2

OpenStudy (anonymous):

y is 4 not 2 y=6-x y=6-2 y=4

OpenStudy (maheshmeghwal9):

You can take any eq. of those two for finding Y.Since u will put the value of x which is equal to 2.

OpenStudy (anonymous):

Yes but y=-x+6=-2+4=2 is not correct it should read y=-x+6=-2+6=4

OpenStudy (anonymous):

y=5x-6 y=5*2-6 y=10-6 y=4

OpenStudy (anonymous):

x=2 and y=4 is the solution

OpenStudy (maheshmeghwal9):

Oops!Sorry,but you got it.That's the answer.

OpenStudy (anonymous):

Im not asking the question im answering it

OpenStudy (maheshmeghwal9):

Ya,I know but think once again upon my saying.